40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Oct 15 08:12:45 2012, Steve, Update, VAC, blank vac monitor blankvac.png
    Reply  Mon Oct 15 13:07:51 2012, jamie, Update, VAC, vacuum VME crate broken, replaced, minor vacuum mayhem ensues 
       Reply  Mon Oct 15 14:51:16 2012, Jenne, Update, SUS, SUS were kicked hard as a result 
          Reply  Mon Oct 15 17:08:01 2012, Steve, Update, VAC, vacuum is back with some disabled valves VMEcrateout.pngIMG_1725.JPGwarning.pngshortcomings.jpg
             Reply  Wed Oct 24 09:15:12 2012, Steve, Update, VAC, vacuum rack 24V ps was current limited currentlimited24V.jpg
                Reply  Mon Nov 5 10:34:48 2012, Steve, Update, VAC, existing vacuum documents 
                Reply  Fri Aug 16 09:35:18 2013, Steve, Update, VAC, Vacuum Normal state recognition is back VacuumNormal.pngvacValvesDisabled.jpg
Message ID: 9017     Entry time: Fri Aug 16 09:35:18 2013     In reply to: 7605
Author: Steve 
Type: Update 
Category: VAC 
Subject: Vacuum Normal state recognition is back 

Quote:

Quote:

Quote:

Quote:

Apparently all of the ION pump valves (VIPEE, VIPEV, VIPSV, VIPSE) opened, which vented the main volume up to 62 mTorr.  All of the annulus valves (VAVSE, VAVSV, VAVBS, VAVEV, VAVEE) also appeared to be open.  One of the roughing pumps was also turned on.  Other stuff we didn't notice?  Bad. 

 Several of the suspensions were kicked pretty hard (600+ mV on some sensors) as a result of this quick vent wind.  All of the suspensions are damped now, so it doesn't look like we suffered any damage to suspensions.

CLOSE CALL on the vacuum system:

Jamie and I disabled V1, VM2 and VM3 gate valves by disconnecting their 120V solenoid actuator before the swap of the VME crate.

The vacuum controller unexpectedly lost control over the swap as Jamie described it. We were lucky not to do any damage! The ion pumps were cold and clean. We have not used them for years so their outgassing possibly  accumulated to reach ~10-50 Torr

I disconnected_ immobilized and labelled the following 6 valves:  the 4 large ion pump gate valves and VC1,  VC2  of the cryo pump. Note: the valves on the cryo pump stayed closed. It is crucial that a warm cry pump is kept closed!

This will not allow the same thing to happen again and protect the IFO from warm cryo contamination.

The down side of this that the computer can not identify vacuum states any longer.

This vacuum system badly needs an upgrade. I will make a list.

 While I was doing the oil change of the roughing pumps I accidentally touched the 24 V adjustment knob on the power supply.

All valve closed to default condition. I realized that the current indicator was red at 0.2A  and the voltage fluctuated from 3-13V

Increased current limiter to 0.4A and set voltage to 24V     I think this was the reason for the caos of valve switching during the VME swap.

 

 Based on the facts above I reconnected VC1 and VC2 valves.  State recognition is working.  Ion pumps are turned off and their gate valves are disabled. 

We learned that even with closed off gate valves while at atmosphere  ion pumps outgass hydrocarbons at 1e-6 Torr level.  We have not used them for this reason in the passed 9 rears.

 

I need help with implementing V1 interlock triggered by Maglev failure signal  and-or P2 pressure.

MEDM screen agrees with vacuum rack signs.

Attachment 1: VacuumNormal.png  393 kB  | Hide | Hide all
VacuumNormal.png
Attachment 2: vacValvesDisabled.jpg  144 kB  Uploaded Fri Aug 16 11:05:06 2013  | Hide | Hide all
vacValvesDisabled.jpg
ELOG V3.1.3-